OpenStudy (anonymous):

The area of a square garden is 202 feet. Estimate the side length of the garden.

OpenStudy (anonymous):

area of a square = side * side = side^2 now given that the area of the square garden is 202 sq. feet means that side^2 = 202 sq. feet side = square root of 202 sq.feet but mainly the area given should be square feet ( in units ) please note that as the area can not be in unit feet it should be in square feet side = 14.2 feet ( approx.)

OpenStudy (anonymous):

\[area= s^2\] \[202=s^2\] \[s=\sqrt{202}\] \[s=14.21\]

OpenStudy (anonymous):

if i did mistake then sorry

OpenStudy (anonymous):

have u got the correct answer basket ball

OpenStudy (anonymous):

mathpro has it

OpenStudy (anonymous):

congrats satelite for becoming as a legend

OpenStudy (anonymous):

thanks.. what anout this problem : -6m - 6 +8m = -5 +2m - 1. Tell whether the equation has infinitely many solutions or no solutions.

OpenStudy (anonymous):

ok see here , -6m + 8m - 6 = -5 + 2m - 1 2m - 6 = - 5 -1 + 2m 2m - 6 = -6 + 2m transpose 2m to LHS and -6 to RHS 2m - 2m = -6 + 6 0 = 0 infinite solutions

OpenStudy (anonymous):

whats LHS & RHS

OpenStudy (anonymous):

left hand side and right hand side

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The coordinates of three vertices of a rectangle are A(1,-5), B(1,3) and C(10,3). find the coordinates of the fourth vertex. then, find the arrea of the rectangle

OpenStudy (anonymous):

it is left hand side and right hand side . ask any more problems

OpenStudy (anonymous):

The coordinates of three vertices of a rectangle are A(1,-5), B(1,3) and C(10,3). find the coordinates of the fourth vertex. then, find the arrea of the rectangle

OpenStudy (anonymous):

The coordinates of three vertices of a rectangle are A(1,-5), B(1,3) and C(10,3). find the coordinates of the fourth vertex. then, find the arrea of the rectangle

OpenStudy (anonymous):

ok , let me answer u

OpenStudy (anonymous):

k

OpenStudy (anonymous):

see the other coordinate D would be ( 10, -5 ) since opposite sides of a rectangle are equal hence D = (10,-5 ) now the area of the rectangle would be (AB) * ( BC ) since area of rectangle = length * breadth now AB = 5 unit and BC = 7 unit hence the area = 7unit * 5 unit = 35 ( unit^2) hope u got it .

OpenStudy (anonymous):

see the other coordinate D would be ( 10, -5 ) since opposite sides of a rectangle are equal hence D = (10,-5 ) now the area of the rectangle would be (AB) * ( BC ) since area of rectangle = length * breadth now AB = 5 unit and BC = 7 unit hence the area = 7unit * 5 unit = 35 ( unit^2) hope u got it .

OpenStudy (anonymous):

see the other coordinate D would be ( 10, -5 ) since opposite sides of a rectangle are equal hence D = (10,-5 ) now the area of the rectangle would be (AB) * ( BC ) since area of rectangle = length * breadth now AB = 5 unit and BC = 7 unit hence the area = 7unit * 5 unit = 35 ( unit^2) hope u got it .

OpenStudy (anonymous):

got it or not

OpenStudy (anonymous):

isnt the area 72u^2

OpenStudy (anonymous):

don't think so but may be . but according to me 35 u^2 is correct one

OpenStudy (anonymous):

k

OpenStudy (anonymous):

The formula for the resistance of a conductor with voltage V and current I is r= V/I. slove for V

OpenStudy (anonymous):

would the answer be R.I=V......?

OpenStudy (anonymous):

yes of course , i think so

OpenStudy (anonymous):

k

OpenStudy (anonymous):

any more questions

OpenStudy (anonymous):

slove -0.25 + 1.75x < -1.75 + 2.25x

OpenStudy (anonymous):

k so there is no equal sign

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

is : 3 < x the answer?

OpenStudy (anonymous):

.....?

OpenStudy (anonymous):

let them be converted into fractions : (-25/100 ) + (175/100)x < (-175/100) + (225/100)x 25(-5/100 + 7/100 x) < 25(-7/100 + 5/100)x 25(-5+7x)/100 < 25(-7+5x)/100 (-5+7x)/100 <( -7 + 5x)/100 -5+7x < -7 + 5x 2 < -2x 2/-2 < x -1 < x actually i don't know how to solve these type of problems but i found this

OpenStudy (anonymous):

oh sorry u r correct

OpenStudy (anonymous):

-0.25+1.75x<-1.75+2.25x Since 2.25x contains the variable to solve for, move it to the left-hand side of the inequality by subtracting 2.25x from both sides. 1.75x-0.25-2.25x<-1.75 Since 1.75x and -2.25x are like terms, add -2.25x to 1.75x to get -0.5x. -0.5x-0.25<-1.75 Since -0.25 does not contain the variable to solve for, move it to the right-hand side of the inequality by adding 0.25 to both sides. -0.5x<0.25-1.75 Subtract 1.75 from 0.25 to get -1.5. -0.5x<-1.5 Divide each term in the inequality by -0.5. -(0.5x)/(-0.5)>-(1.5)/(-0.5) Cancel the common factor of -0.5. x>-(1.5)/(-0.5) Simplify the right-hand side of the inequality by simplifying each term. x>3

OpenStudy (anonymous):

got it now

OpenStudy (anonymous):

-0.25+1.75x<-1.75+2.25x Since 2.25x contains the variable to solve for, move it to the left-hand side of the inequality by subtracting 2.25x from both sides. 1.75x-0.25-2.25x<-1.75 Since 1.75x and -2.25x are like terms, add -2.25x to 1.75x to get -0.5x. -0.5x-0.25<-1.75 Since -0.25 does not contain the variable to solve for, move it to the right-hand side of the inequality by adding 0.25 to both sides. -0.5x<0.25-1.75 Subtract 1.75 from 0.25 to get -1.5. -0.5x<-1.5 Divide each term in the inequality by -0.5. -(0.5x)/(-0.5)>-(1.5)/(-0.5) Cancel the common factor of -0.5. x>-(1.5)/(-0.5) Simplify the right-hand side of the inequality by simplifying each term. x>3

OpenStudy (anonymous):

-0.25+1.75x<-1.75+2.25x Since 2.25x contains the variable to solve for, move it to the left-hand side of the inequality by subtracting 2.25x from both sides. 1.75x-0.25-2.25x<-1.75 Since 1.75x and -2.25x are like terms, add -2.25x to 1.75x to get -0.5x. -0.5x-0.25<-1.75 Since -0.25 does not contain the variable to solve for, move it to the right-hand side of the inequality by adding 0.25 to both sides. -0.5x<0.25-1.75 Subtract 1.75 from 0.25 to get -1.5. -0.5x<-1.5 Divide each term in the inequality by -0.5. -(0.5x)/(-0.5)>-(1.5)/(-0.5) Cancel the common factor of -0.5. x>-(1.5)/(-0.5) Simplify the right-hand side of the inequality by simplifying each term. x>3

OpenStudy (anonymous):

-0.25+1.75x<-1.75+2.25x Since 2.25x contains the variable to solve for, move it to the left-hand side of the inequality by subtracting 2.25x from both sides. 1.75x-0.25-2.25x<-1.75 Since 1.75x and -2.25x are like terms, add -2.25x to 1.75x to get -0.5x. -0.5x-0.25<-1.75 Since -0.25 does not contain the variable to solve for, move it to the right-hand side of the inequality by adding 0.25 to both sides. -0.5x<0.25-1.75 Subtract 1.75 from 0.25 to get -1.5. -0.5x<-1.5 Divide each term in the inequality by -0.5. -(0.5x)/(-0.5)>-(1.5)/(-0.5) Cancel the common factor of -0.5. x>-(1.5)/(-0.5) Simplify the right-hand side of the inequality by simplifying each term. x>3

OpenStudy (anonymous):

-0.25+1.75x<-1.75+2.25x Since 2.25x contains the variable to solve for, move it to the left-hand side of the inequality by subtracting 2.25x from both sides. 1.75x-0.25-2.25x<-1.75 Since 1.75x and -2.25x are like terms, add -2.25x to 1.75x to get -0.5x. -0.5x-0.25<-1.75 Since -0.25 does not contain the variable to solve for, move it to the right-hand side of the inequality by adding 0.25 to both sides. -0.5x<0.25-1.75 Subtract 1.75 from 0.25 to get -1.5. -0.5x<-1.5 Divide each term in the inequality by -0.5. -(0.5x)/(-0.5)>-(1.5)/(-0.5) Cancel the common factor of -0.5. x>-(1.5)/(-0.5) Simplify the right-hand side of the inequality by simplifying each term. x>3

OpenStudy (anonymous):

-0.25+1.75x<-1.75+2.25x Since 2.25x contains the variable to solve for, move it to the left-hand side of the inequality by subtracting 2.25x from both sides. 1.75x-0.25-2.25x<-1.75 Since 1.75x and -2.25x are like terms, add -2.25x to 1.75x to get -0.5x. -0.5x-0.25<-1.75 Since -0.25 does not contain the variable to solve for, move it to the right-hand side of the inequality by adding 0.25 to both sides. -0.5x<0.25-1.75 Subtract 1.75 from 0.25 to get -1.5. -0.5x<-1.5 Divide each term in the inequality by -0.5. -(0.5x)/(-0.5)>-(1.5)/(-0.5) Cancel the common factor of -0.5. x>-(1.5)/(-0.5) Simplify the right-hand side of the inequality by simplifying each term. x>3